# G-Adaptivity: optimised graph-based mesh relocation for finite element methods

### v2

@@ Appendix F Mesh Tangling Prevention
(Mesh Tangling). A mesh is said to be tangled if at least one simplex in the triangulation has a negative determinant in its Jacobian matrix, i.e., $\det(J_{i})\leq 0,\quad\text{for some }i$, where $J_{i}$ is the Jacobian matrix of the affine transformation mapping the reference element to the physical element in the mesh. Equivalently, the mesh is untangled if all eigenvalues of the Hessian of the transformation function, or its discrete counterpart given by the graph Laplacian, remain positive.
@@ F.1 Continuous-Time Evolution
Since $\operatorname{tr}(\Delta)=N\geq 0$, we have $J(t)>0$ for all $t\geq 0$, ensuring that no elements invert.
@@ F.2 Time Step Constraints for Mesh Preservation
Given the discrete update $\textbf{X}^{k+1}=(I-dt\Delta)\textbf{X}^{k}$, the mesh remains untangled if $dt\leq\frac{1}{2}$.
@@
The determinant of the deformation Jacobian propagates as $\det J^{k+1}=\det(I-dt\Delta)\det J^{k}$. To ensure $\det J^{k+1}>0$, we require $\det(I-dt\Delta)>0$. The eigenvalues of $I-dt\Delta$ are $\mu_{i}=1-dt\lambda_{i}$ with $0\leq\lambda_{i}\leq 2$, so the determinant condition reduces to
@@
Ensuring positivity requires $01-dt\lambda_{\max}>$ 0. Now by applying Gershgorin’s theorem to $A-I$ and noting that $A$ has positive entries with row sum equal to 1, it follows that $\lambda_{\max}=2$. Hence we obtain $dt<\frac{1}{2}$. ∎
@@ F.4 Mesh Quality Measures
Mesh quality measures are often used as indicators of whether a mesh will be effective when used to solve a PDE. As our method is designed to minimise the FE solution error directly the importance of such quality measures is reduced for our mesh generation algorithm, however they still serve as useful regularizers for the loss function used by the GNN to construct the mesh. These quality measures assess scale, skewness, and regularity, ensuring a well-conditioned mesh.

### v3

@@ Appendix F Mesh Tangling Prevention
(Mesh Tangling). We say that a physical mesh is tangled if at least one simplex in the triangulation has a negative determinant in its Jacobian matrix, i.e., $\det(J_{i})\leq 0,\,\text{for some }i$, where $J_{i}$ is the Jacobian matrix of the affine transformation mapping the reference element to the physical element in the mesh. Equivalently, the mesh is untangled if all eigenvalues of the Hessian of the transformation function, or its discrete counterpart given by the graph Laplacian, remain positive.
@@ F.1 Continuous-Time Evolution
Since $\operatorname{tr}(\Delta)={N_{x}\times d}\geq 0$, we have $J(t)>0$ for all $t\geq 0$, ensuring that no elements invert.
@@ F.2 Time Step Constraints for Mesh Preservation
Given the discrete update $\textbf{X}^{k+1}=(I-dt\Delta)\textbf{X}^{k}$, the mesh remains untangled if $dt<\frac{1}{2}$.
@@
The determinant of the deformation Jacobian propagates as $\det J^{k+1}=\det(I-dt\Delta)\det J^{k}$. To ensure $\det J^{k+1}>0$, we require $\det(I-dt\Delta)>0$. The eigenvalues of $I-dt\Delta$ are $\mu_{i}=1-dt\lambda_{i}$ so the determinant condition reduces to
@@
Noting that $A$ has positive entries with row sum equal to 1, it follows by Gershgorin’s theorem that the eigenvalues of $A-I$ are contained in the Gershgorin circle $|\lambda_{i}-1|<1$. Seeing as the coefficients of $A-I$ are real-valued the eigenvalues of $A-I$ are either real-valued or come in complex conjugate pairs. If $\lambda_{i}$ is real-valued the contribution to the above determinant is $01-dt\lambda_{i}>1-2dt>$ 0 if $dt>1/2$. If $\mathrm{Im}\lambda_{i}\neq 0$ then $\overline{\lambda_{i}}$ is also an eigenvalue and the contribution to the determinant is $(1-dt\lambda_{i})(1-dt\overline{\lambda_{i}})=|1-dt\mathrm{Re}\lambda_{i}|^{2}+|\mathrm{Im}\lambda_{i}|^{2}>0$. Hence we obtain $\det(I-dt\Delta)>0$ if $dt<\frac{1}{2}$. ∎
@@ F.4 Mesh Quality Measures
Mesh quality measures are often used as indicators of whether a mesh will be effective when used to solve a PDE. In particular, the mesh-quality metrics are directly related to the conditioning of FEM stiffness matrices, meaning poor mesh conditioning leads to numerical instabilities in the FEM solvers. Our method is designed to minimise the FE solution error directly, but the inclusion of the equidistribution regularisation ensures that G-Adaptivity leads to meshes that maintain good conditioning while reducing the FEM error. In our numerical experiments we report aspect ratio as a strong indicator of this mesh conditioning, but in the relevant literature the following metrics are commonly used to often used to assess mesh scale, skewness, and regularity.